OpenStudy (anonymous):

State the restrictions and simplify: 1/x^2 + 1/x-2

OpenStudy (anonymous):

\[\frac{ 1 }{ ^{x2} } + \frac{ 1 }{ x-2 }\]

OpenStudy (allank):

Restriction: x must not =2.

OpenStudy (allank):

To simplify, begin by combining the two terms. Hint: use the least common multiple as the denominator.

OpenStudy (anonymous):

@jim_thompson5910 got another for ya

OpenStudy (anonymous):



OpenStudy (anonymous):

@allank

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

is it x-2 or x+2?

OpenStudy (anonymous):

x-2

OpenStudy (anonymous):

sorry my drawing I put x+2 in error

OpenStudy (allank):

First, what do you think is the limitation for this?

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

\[\Large \frac{ 1 }{ x^{2} } + \frac{ 1 }{ x-2 }\] \[\Large \frac{ 1(x-2) }{ x^{2}(x-2) } + \frac{ 1 }{ x-2 }\] \[\Large \frac{ x-2 }{ x^{2}(x-2) } + \frac{ 1 }{ x-2 }\] \[\Large \frac{ x-2 }{ x^{2}(x-2) } + \frac{ 1*x^{2} }{ x^{2}(x-2) }\] \[\Large \frac{ x-2 }{ x^{2}(x-2) } + \frac{ x^{2} }{ x^{2}(x-2) }\] \[\Large \frac{ x-2 + x^{2} }{ x^{2}(x-2) }\] \[\Large \frac{ x^{2} + x-2 }{ x^{2}(x-2) }\] =================================================================================== So \[\Large \frac{ 1 }{ x^{2} } + \frac{ 1 }{ x-2 }\] simplifies to \[\Large \frac{ x^{2} + x-2 }{ x^{2}(x-2) }\]

jimthompson5910 (jim_thompson5910):

We're following the same idea as last time. The big difference this time we're adding instead of subtracting the fractions.

OpenStudy (anonymous):

@jim_thompson5910 that is what I had but I was not seeing the bigger picture. I needed to rearrange it for x^2 to x to -2 so I can then factor it out. Ugg

jimthompson5910 (jim_thompson5910):

you can factor x^2 + x - 2 to get (x+2)(x-1), but this is optional in my opinion

OpenStudy (anonymous):

@jim_thompson5910 I really appreciate your help

jimthompson5910 (jim_thompson5910):

yw, glad to be of help again